OpenStudy (anonymous):

Maggie's brother is three years younger than twice her age. The sum of their ages is 24. How old is maggie

OpenStudy (ashes.boom):

maggies brother = b maggie = m b = 3 - 2m

OpenStudy (anonymous):

How would I solve it

OpenStudy (ashes.boom):

then it says their sum is equal to 24 so you got b+m = 24

OpenStudy (ashes.boom):

then you can substitute the b for (3-2m) and you got m + (3-2m) = 24 and you can solve for m

OpenStudy (ashes.boom):

Make sense?

OpenStudy (anonymous):

Then would u combined like turms or distribute

OpenStudy (ashes.boom):

i think you distribute

OpenStudy (zenshiru):

Essentially, what @Ashes.boom said is correct, but I'll just word it a different way. Let's put it like this. M = the age of Maggie. B = the age of her brother. Maggie's brother is 3 years younger than twice her age, so we'll get this formula: b = 2m - 3 Now, it's said the sum of their ages is 24, so you now have: m + b = 24 From here, you'll replace b with (2m - 3), and get this: m + (2m - 3) = 24 Now you can combine 2m and m, which gives you 3m. So you're now left with: 3m - 3 = 24 Now, do a reverse operation and add the three to both sides, and it'll cancel on the left side, so you now have: 3m = 27 Now just divide by 3, and that'll tell you Maggie's age.

OpenStudy (anonymous):

Solve the following for b and m, Maggie's brother's age and Maggie's age. \[\{b=2 m-3,b+m=24\} \]